OpenStudy (anonymous):

Please help! I'll give a medal. Let vec a = -4 i-2 j+1 k and vec b = 0 i-5 j+5 k. Find the angle between the vectors

OpenStudy (dan815):

dot product a . b = |a| |b| cos theta

OpenStudy (dan815):

theta is the angle between these 2 vectors

OpenStudy (anonymous):

So it would be cos theta=\[\frac{ 15 }{ \sqrt{21}\sqrt{50} }\] which results in theta being about 62.42?

OpenStudy (dan815):

yes